OpenStudy (anonymous):

Find the vertex of the following function: f (x) = 3x^2 + 6x - 5 .

OpenStudy (anonymous):

the formula is: -b/2a ?

OpenStudy (anonymous):

x of vertex = -b/2a y of vertex, insert x of vertex into standard formula and solve for y

OpenStudy (anonymous):

I got -b/2a = -(6/2(3) = -(6/6) = -1 ?

OpenStudy (anonymous):

let me work it and check for you

jimthompson5910 (jim_thompson5910):

mocham, you have the correct x coordinate of the vertex

OpenStudy (anonymous):

yes, now what's the y?

OpenStudy (anonymous):

the question doesn't specify if it wants the x or y vertex BUT I have two possible answers so it probably wants it - that makes sense

jimthompson5910 (jim_thompson5910):

yes you need both since the vertex is in the form (x,y)

jimthompson5910 (jim_thompson5910):

ie it's a point

OpenStudy (anonymous):

normally it will be in the form of (x,y) sometimes they will ask the height of the vertex which is y sometimes they will ask the location of the vertex which will be x

OpenStudy (anonymous):

depends on the wording of the question, beware of trickery

OpenStudy (anonymous):

now its k=f(-b/2a) ?

OpenStudy (anonymous):

f(-b/2a) like f(-2) where yhey want you to evaluate function at x = -2 k=f(-b/2a) should mean that they want the y of the vertex. Anyone concur with this statement?

OpenStudy (anonymous):

the statement is asking you to evaluate the function at the x of the vertex

OpenStudy (anonymous):

and return it's y value

OpenStudy (anonymous):

k= -1(-b/2a) wouldn't my answer be 1?

jimthompson5910 (jim_thompson5910):

mocham, you found above that -b/2a = -1 so f(-b/2a) = f(-1) = ???

OpenStudy (anonymous):

-1?

jimthompson5910 (jim_thompson5910):

f (x) = 3x^2 + 6x - 5 f(-1) = 3(-1)^2 + 6(-1) - 5 f(-1) = ???

OpenStudy (anonymous):

k= f(-b/2a) k= f(-1) y = f(-1) all of these are asking the same thing

OpenStudy (anonymous):